A 100W, 250V lamp is connected in parallel with an unknown resistance ‘R,’ across a 250 V supply. The total power dissipated in the circuit is 1100V. Find the value of unknown resistance. Assume the resistance of the lamp remains unaltered. 


1
Expert's answer
2021-02-25T11:23:41-0500

Answer

Resistance of given lamp

"R=\frac{V^2}{P}=\frac{250\times250}{100}=625\Omega"


Now let unknown resistance is R

So equivalent of these resistance

"R_eq=\frac{625R}{625+R}"

Now total power

"P=VI=\frac{V^2}{R}"

"1100=\frac{250\times250(625+R)}{625R}"

Now solute

"11R=625+R"

"R=56.8\Omega"